# Finite Expression Method for Learning Dynamics on Complex Networks

### v1

@@ 3.4 Implementation of FEX
As mentioned previously, FEX’s computational process begins with the formation of a mathematical binary tree, where each node represents either a unary or a binary operator. The candidate solution $u$ is derived from evaluating the function represented by this tree. Subsequently, the CO denoted by  (3) is employed to dynamically choose the best operators for all tree nodes. This optimization aims to find operators that recover the structure of the genuine solution.
@@ 3.4.1 Finite expressions with binary trees
FEX employs a binary tree structure $\mathcal{T}$ to represent finite expressions. The user defines the sets of potential unary and binary operators, represented by $\mathbb{U}$ and $\mathbb{B}$. Common choices of unary and binary operators are respectively,
@@ 3.4.2 Implementation of FEX
  • Score Computation: To discover the solution’s structure, FEX deploys a mix-order optimization algorithm. This step assesses the score of the operator sequence $\bm{e}$. RBM is employed to accelerate the score’s calculation over the network data.
@@ 3.4.3 Score computation
Random Batch Method To address the computational burden when performing the total $T_{1}+T_{2}$ optimization steps, we propose to approximate the interaction terms using only a subset of the particles at each step. In particular, we divide the $N$ particles into $n$ batches and only consider the interactions within each batch, then the number of interactions that need to be computed is significantly reduced. Assuming that each batch contains approximately $p$ nodes ($n\approx\frac{N}{p}$), then the number of interactions to be computed for each batch is on the order of $p^{2}$. Since there are $n$ such batches, the total number of interactions at each step is on the order of $np^{2}=\frac{N}{p}p^{2}=Np$, which is $O(N)$ since $p$ is a constant independent of $N$. This significantly reduces the computational cost when evaluating the loss function during the $T_{1}+T_{2}$ steps of optimization.
@@ 3.4.5 Controller Update
where $\mathbf{1}$ denotes an indicator function which evaluates to $1$ when the condition is met, and 00 otherwise. Meanwhile, $\hat{S}_{\nu,\Phi}$ represents the $(1-\nu)$-quantile of the scores within the set $\{S(\mathbf{e}^{(i)})\}_{i=1}^{N}$.
@@ 4 Dynamics learning results
In this section, we present the results of identification of HR, FHN, and coupled Rössler dynamics with FEX on the Scale-Free (SF) network with clean data. The set up of SF network follows the description in Appendix 6.1. We choose SF network since it demonstrates more complex phenomena than the ER model, and many real-world networks have scale-free property. During the offline stage, we generate the time series data with the true dynamics (ODEs). During the inference stage, we use FEX only with the time series data to identify the governing equations of the system, without accessing to the genuine ODEs. Throughout all numerical experiments, we set the batch size in RBM as 32. In addition, we provide the depths of FEX trees used in each of the considered dynamics in the numerical examples in Table 1.
@@ 4.2 Hindmarsh-Rose dynamics
[…]
@@ 4.3 FitzHugh-Nagumo dynamics
where the first component $x_{i,1}$ represents the membrane potential containing self and interaction dynamics, $k_{i}^{in}$ is the in-degree of neuron $i$ (representing the number of incoming connections to node $i$), and $\epsilon=1$. The second component $x_{i,2}$ represents a recovery variable where $a=0.28,b=0.5$ and $c=-0.04$. We generate the time series data with Equation (16), with the adjacency matrix of SF network, terminal time $T=300$ and time step $\delta t=0.01$.
@@ 4.4 Coupled Rössler dynamics
where coupling strength $\epsilon=0.15$, $a=0.2,b=0.2,$ and $c=-5.7$. The natural frequencies of the oscillators, denoted by $\omega_{i}$, follow a normal distribution with a mean value of 1 and a standard deviation of 0.1. We generate the time series data with Equation (17), with the adjacency matrix of SF network, terminal time $T=100$ and time step $\delta t=0.01$.
@@
4.5 Details of performance of random batch method
As introduced before, computation on a large complex network is demanding or even infeasible. Therefore, we incorporate the RBM strategy into the FEX algorithm to accelerate the identification of the optimal mathematical structure of network dynamics while maintaining a high accuracy.
To demonstrate the efficiency of RBM in accelerating CO in the training process of FEX, we create a ER model of network size $64,128,256$, and $512$, respectively. Then, we use the time series data of FHN model on the four ER models, and implement FEX to infer the dynamics. We record the clock time of computation of $T_{1}+T_{2}$ coarse-tune in the FEX training algorithm by full interaction method and RBM (with a batch size of 32), respectively. The log-log plot of time vs. number of particles is shown in Fig. 7. We observe that the fitting line clock time of the full interaction model on different network sizes has a slope of approximately 2. The RBM’s slope is roughly 1, indicating that the RBM demonstrates a linear scaling performance.
Figure 7: Clock time comparisons of full interaction model and RBM. The $x$-axis shows 64, 128, 256, and 512 particles of the ER network, respectively. The $y$-axis shows the clock time of one iteration computation for each network size.
Furthermore, we comment that we only use RBM strategy in the CO part of FEX to identify the optimal structure of the dynamics. The CO stage is the computational bottleneck of FEX, and RBM has demonstrated its capability to determine the correct structure of dynamics while accelerating the computation significantly. After the structures of top-performing candidates are learned, we use the full interaction model in the $T_{3}$ iterations of coefficients fine-tuning stage to achieve high accuracy of dynamics inferred by FEX. The computational burden at this stage is manageable since we only have to fine-tune the coefficients of the top $K$ candidates in the pool $\mathbb{P}$.
@@ 5 Conclusion
In this work, we proposed FEX as a novel methodology to learn the dynamics on complex networks. FEX has shown its capability to accurately discover physical laws on various synthetic networks, even with noise and low resolution of data. Furthermore, we incorporate RBM into our proposed FEX algorithm, which scales up FEX to handle large data when the network size increases significantly. Therefore, we have demonstrated that FEX has the potential to identify real-world dynamics in an interpretable and reliable way.

### v3

@@ 3.4 Implementation of FEX
As mentioned previously, FEX’s computational process begins with the formation of a mathematical binary tree, where each node represents either a unary or a binary operator. The candidate solution is derived from evaluating the function represented by this tree. Subsequently, the CO denoted by  (3) is employed to dynamically choose the best operators for all tree nodes. This optimization aims to find operators that recover the structure of the genuine network dynamics.
@@ 3.4.1 Finite expressions with binary trees
FEX uses a binary tree structure $\mathcal{T}$ to represent finite expressions. The user defines the sets of potential unary and binary operators, represented by $\mathbb{U}$ and $\mathbb{B}$. Common choices of unary and binary operators are respectively,
@@ 3.4.2 Workflow of FEX
  • Score Computation: To discover the solution’s structure, FEX deploys a mix-order optimization algorithm. This step assesses the score of the operator sequence $\bm{e}$. A stochastic algorithm is utilized to accelerate the score’s calculation over the network data.
@@ 3.4.3 Score computation
Stochastic-FEX To address the computational burden when performing the total $T_{1}+T_{2}$ optimization steps, we propose to approximate the interaction terms using only a subset of the particles at each step. In particular, we divide the $N$ particles into $n$ batches and only consider the interactions within each batch, then the number of interactions that need to be computed is significantly reduced. Assuming that each batch contains approximately $p$ nodes ($n\approx\frac{N}{p}$), then the number of interactions to be computed for each batch is on the order of $p^{2}$. Since there are $n$ such batches, the total number of interactions at each step is on the order of $np^{2}=\frac{N}{p}p^{2}=Np$, which is $O(N)$ since $p$ is a constant independent of $N$. This significantly reduces the computational cost when evaluating the loss function during the $T_{1}+T_{2}$ steps of optimization.
@@
We have the following theorem on the convergence of Stochastic-FEX within this phase. The proof is provided in the Supplementary Information.
Theorem 1.
Fix some tolerance level $\delta>0$, and integer $q\geq 2$. Let $\theta^{*}=(\theta_{F}^{*},\theta_{G}^{*})$ be a regular minimizer of $\mathcal{L}$ defined in (2), and suppose that Stochastic-FEX is run with a step-size schedule of the form $\gamma_{n}=\gamma/(n+m)^{p}$ for some $p\in(2/(q+2),1]$ and large enough $m,\gamma>0$. Furthermore, assume that the following bounds hold:
  1. 1.
    $\left\|\frac{d\mathbf{x}_{i}(t)}{dt}-\mathbf{F}(\mathbf{x}_{i}(t);\theta_{F})\right\|_{2}\leq C_{1}$
  2. 2.
    $\left\|\mathbf{G}(\mathbf{x}_{i}(t),\mathbf{x}_{j}(t);\theta_{G})\right\|_{2}\leq M_{1}$
  3. 3.
    $\left|\nabla_{\theta_{G}}\mathbf{G}(\mathbf{x}_{i}(t),\mathbf{x}_{j}(t);\theta_{G})_{k}\right|\leq M_{2}$, where $(\cdot)_{k}$ denotes the $k$-th component of the vector.
Then:
  1. 1.
    There exist neighborhoods $\mathcal{U}$ and $\mathcal{U}_{1}$ of $\theta^{*}$ such that, if $\theta_{1}\in\mathcal{U}_{1}$, the event
    $\Omega_{\mathcal{U}}=\{\theta_{n}\in\mathcal{U}\text{ for all }n=1,2,\ldots\}$
    occurs with probability at least $1-\delta$.
  2. 2.
    Conditioned on $\Omega_{\mathcal{U}}$, we have
    $\mathbb{E}[\|\theta_{n}-\theta^{*}\|_{2}^{2}\mid\Omega_{\mathcal{U}}]=O(1/n^{p}).$
@@ 3.4.5 Controller update
where $\mathbf{1}$ denotes an indicator function which evaluates to $1$ when the condition is met, and 0 otherwise. Meanwhile, $\hat{S}_{\nu,\Phi}$ represents the $(1-\nu)$-quantile of the scores within the set $\{S(\mathbf{e}^{(i)})\}_{i=1}^{M}$.
@@ 4 Dynamics learning results
In this section, we present the results of identification of HR, FHN, and coupled Rössler dynamics with FEX on the Scale-Free (SF) network with clean data. The set up of SF network follows the description in Supplementary Information. We choose SF network since it demonstrates more complex phenomena than the ER model, and many real-world networks have scale-free property. During the offline stage, we generate the time series data with the true dynamics (ODEs). During the inference stage, we use FEX only with the time series data to identify the governing equations of the system, without accessing to the genuine ODEs. We emphasize that we only have access to the time series data and approximate the state time derivatives using numerical derivatives. This contrasts with the SINDy method, which requires direct access to the time derivatives of the states. Throughout all numerical experiments, we set the batch size in stochastic-FEX as 32. In addition, we provide the depths of FEX trees used in each of the considered dynamics in the numerical examples in Table 1. Robustness analyses of FEX and other baseline methods, aimed at extracting dynamics from low-resolution data, networks with spurious or missing links, and noisy data, are discussed in the Supplementary Information.
@@ 4.2 Hindmarsh-Rose dynamics
$\left\{\begin{aligned} \frac{d\hat{x}_{i,1}}{dt}&=0.9961x_{i,2}-1.0014x_{i,1}^{3}+2.9758x_{i,1}^{2}-0.9917x_{i,3}+3.2392\\ &+\sum_{j=1}^{N}\mathbf{A}_{ij}\left(0.2985\sigma\left(x_{j,1}\right)-0.1477x_{i,1}\sigma\left(x_{j,1}\right)\right),\\ \frac{d\hat{x}_{i,2}}{dt}&=1.0004-5.0001x_{i,1}^{2}-1.0001x_{i,2},\\ \frac{d\hat{x}_{i,3}}{dt}&=0.0255+0.0182x_{i,1}-0.0050x_{i,3},\end{aligned}\right$. (15)
@@
4.2.1 Observational noises
In this subsection, we demonstrate the robustness of our method against observational noises. We focus on the HR dynamics on a directed SF network. We generate noisy data through
$X_{i}^{obs}(t)=\bm{x}_{i}(t)+a\beta_{X}(t),$
where $\beta_{X}(t)$ is observational noise, which follows a Gaussian distribution with zero mean and standard deviation one, so the parameter $a$ is used to tune the level of noise.
True Dynamics Two-Phase ARNI FEX
$\mathbf{F}(x_{i})$ $x_{i,2}$ $0.9631x_{i,2}$ $0.6002x_{i,2}$ $0.9916x_{i,2}$
$x_{i,1}^{3}$ $1.0039x_{i,1}^{3}$ $1.3202x_{i,1}^{3}$ $1.0023x_{i,1}^{3}$
$3x_{i,1}^{2}$ $2.7820x_{i,1}^{2}$ $2.3557x_{i,1}^{2}$ $2.9659x_{i,1}^{2}$
$x_{i,3}$ $0.9663x_{i,3}$ $0.7982x_{i,3}$ $0.9871x_{i,3}$
$3.24$ $2.9673$ $1.9699$ $3.2453$
$0.6597\exp(x_{i,1})$
$\mathbf{G}(x_{i},x_{j})$ $0.3\sigma(x_{j,1})$ $0.1529\sigma(x_{j,1})$ $0.2804\sigma(x_{j,1})$ $0.2963\sigma(x_{j,1})$
[…]
Table 2: Numerical results for Hindmarsh-Rose dynamics (SNR = 45) in self-dynamics $\mathbf{F}(x_{i})$ and interaction dynamics $\mathbf{G}(x_{i},x_{j})$ term by term.
True Dynamics Two-Phase ARNI FEX
$\mathbf{F}(x_{i})$ $x_{i,2}$ $0.9341x_{i,2}$ $0.3916x_{i,2}$ $0.9633x_{i,2}$
$x_{i,1}^{3}$ $0.9948x_{i,1}^{3}$ $1.0638x_{i,1}^{3}$ $1.0058x_{i,1}^{3}$
$3x_{i,1}^{2}$ $2.6599x_{i,1}^{2}$ $2.5882x_{i,1}^{2}$ $2.8923x_{i,1}^{2}$
$x_{i,3}$ $0.9332x_{i,3}$ $0.3871x_{i,3}$ $0.9588x_{i,3}$
$3.24$ $3.1341$ $1.8715$ $3.2342$
$0.1318x_{i,2}x_{i,3}$
$0.2246\exp(x_{i,2})$
$\mathbf{G}(x_{i},x_{j})$ $0.3\sigma(x_{j,1})$ $0.1333\sigma(x_{j,1})$ None $0.2683\sigma(x_{j,1})$
$0.15x_{i,1}\sigma(x_{j,1})$ None None $0.1118x_{i,1}\sigma(x_{j,1})$
$0.2341\sigma(x_{j,1}-x_{i,1})$
Table 3: Numerical results for Hindmarsh-Rose dynamics (SNR = 40) in self-dynamics $\mathbf{F}(x_{i})$ and interaction dynamics $\mathbf{G}(x_{i},x_{j})$ term by term.
SNR Two-Phase ARNI FEX
$45$ 0.0924 0.2720 0.0158
$40$ 0.2190 0.6252 0.0376
Table 4: sMAPE of HR dynamics with different levels of noise.
From Table 2, 3 and 4, FEX demonstrates its capability to accurately identify all terms within the HR dynamics with their respective coefficients with SNR = 45 and 40. The Two-Phase method can indeed infer the correct terms in the dynamics in the case of SNR = 45. However, the accuracy of inferred coefficients is low, particularly in the interaction dynamics. In addition, in the case of SNR = 40, Two-Phase method fails to infer the $0.15x_{i,1}\sigma(x_{j,1})$ term in the interaction dynamics. ARNI can infer most of the terms correctly when SNR = 45, with an extra term $0.6597\exp(x_{i,1})$ in the self-dynamics. Furthermore, the coefficients in the self-dynamics are poorly estimated, but it is interesting to notice that the coefficients of interaction-dynamics are more accurate than those of Two-Phase, owing again to ARNI’s capability to infer the network structure. As for SNR = 40, ARNI infers more redundant terms in self-dynamics, and it fails to identify both true terms in the original dynamics, and incorrectly identifies a $0.2341\sigma(x_{j,1}-x_{i,1})$ term. This shows that ARNI can reasonably well infer most of the dynamics terms in low-noise case, while it fails to work as noise level increases. Therefore, FEX has demonstrated its superior performance compared to Two-Phase and ARNI to handle noisy data.
@@ 4.3 FitzHugh-Nagumo dynamics
where the first component $x_{i,1}$ represents the membrane potential containing self and interaction dynamics, $k_{i}^{in}$ is the in-degree of neuron $i$ (representing the number of incoming connections to node $i$), and $\epsilon=1$. The second component $x_{i,2}$ represents a recovery variable where $a=0.28,b=0.5$ and c=0.04c=-0.04. We generate the time series data with Equation (16), with the adjacency matrix of SF network, terminal time $T=300$ and time step $\delta t=0.01$.
@@
We compare the robustness of FEX against spurious and missing links using FHN dynamics, and evaluate it alongside Two-Phase and ARNI methods. FEX demonstrates superior accuracy compared to the baselines, even with random addition or deletion of $15\%$ of the links. Detailed results are provided in the Supplementary Information.
@@ 4.4 Coupled Rössler dynamics
where coupling strength $\epsilon=0.15$, $a=0.2,b=0.2,$ and c=5.7c=-5.7. The natural frequencies of the oscillators, denoted by $\omega_{i}$, follow a normal distribution with a mean value of 1 and a standard deviation of 0.1. We generate the time series data with Equation (17), with the adjacency matrix of SF network, terminal time $T=100$ and time step $\delta t=0.01$.
@@
Utilizing Coupled Rössler dynamics, we further assess the robustness of FEX against low-resolution observational data. We compare the performance of FEX with Two-Phase and ARNI methods when observing only $10\%$ and $5\%$ of the time series data. FEX demonstrates superior accuracy in identifying the dynamics within this low data regime compared to the other baselines. Detailed results are provided in the Supplementary Information.
4.5 Details of performance of stochastic-FEX
As introduced before, computation on a large complex network is demanding or even infeasible. Therefore, we incorporate a stochastic algorithm into the FEX algorithm to accelerate the identification of the optimal mathematical structure of network dynamics while maintaining a high accuracy.
To demonstrate the efficiency, we create a SF model of network size $64,128,256$, and $512$, respectively. Then, we use the time series data of FHN model on the four SF networks, and implement FEX to infer the dynamics. We record the clock time of computation of $T_{1}+T_{2}$ coarse-tune in the FEX training algorithm by full interaction method and stochastic-FEX (with a batch size of 32), respectively. The log-log plot of time vs. number of particles is shown in Fig. 7. We observe that the fitting line clock time of the full interaction model on different network sizes has a slope of approximately 2. The stochastic-FEX’s slope is roughly 1, indicating that the stochastic-FEX demonstrates a linear scaling performance.
Figure 7: Clock time comparisons of full interaction model and stochastic-FEX. The $x$-axis shows 64, 128, 256, and 512 particles of the SF network, respectively. The $y$-axis shows the clock time of one iteration computation for each network size.
Furthermore, we comment that we only use this stochastic algorithm in the CO part of FEX to identify the optimal structure of the dynamics. The CO stage is the computational bottleneck of FEX, and we have demonstrated its capability to determine the correct structure of dynamics while accelerating the computation significantly. After the structures of top-performing candidates are learned, we use the full interaction model in the $T_{3}$ iterations of coefficients fine-tuning stage to achieve high accuracy of dynamics inferred by FEX. The computational burden at this stage is manageable since we only have to fine-tune the coefficients of the top $K$ candidates in the pool $\mathbb{P}$.
@@ 5 Conclusion
In this work, we proposed FEX and its fast algorithm as a novel methodology to learn the dynamics on complex networks. FEX has shown its capability to accurately discover physical laws on various synthetic networks, even with noise and low resolution of data. Furthermore, we incorporate a stochastic algorithm into our proposed FEX framework, which scales up FEX to handle large data when the network size increases significantly. Therefore, we have demonstrated that FEX has the potential to identify real-world dynamics in an interpretable and reliable way.
@@
Acknowledgments C. W. was partially supported by National Science Foundation under award DMS-2206332. H. Y. was partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and DARPA D24AP00325-00.
Competing interests The authors declare no competing interests.